Preferred Stock Valuation

Several years ago, Rolen Riders issued preferred stock with a stated annual dividend of 10% of its $100 par value. Preferred stock of this type currently yields 8%. Assume dividends are paid annually.

  1. What is the estimated value of Rolen's preferred stock? Round your answer to the nearest cent.
    $
  2. Suppose interest rate levels have risen to the point where the preferred stock now yields 13%. What would be the new estimated value of Rolen's preferred stock? Round your answer to the nearest cent.
    $
Answer #1

a.

Dividend = 10% of 100 = 10

Rate of return or yield = 8%

Estimated value of stock = Dividend / Rate of return = 10/8% = 125

b.

New rate of return = 13%,

New estimated value of stock = 10/13% = 76.92